OpenStudy (anonymous):

What is the equation of a circle with center at (-7,-6) and radius of 2??

OpenStudy (anonymous):

the form of a circle: (a-x)^2+(b-y)^2 = r^2 (x, y) = (-7, -6) r = radius

OpenStudy (anonymous):

plug your variables in.

OpenStudy (anonymous):

er, well technically it should be (x-a)2+(y-b)^2 = r^2 (a, b) = (-7, -6) r = radius)

OpenStudy (anonymous):

(x+7)^2+(x+6)^2=4

OpenStudy (agent0smith):

@bloopman your first formula is equivalent to the second hehe. If you factor out a negative from each of the brackets in: \[\Large (a-x)^2+(b-y)^2 = r^2\]you get \[\Large (x-a)^2+(y-b)^2 = r^2\]

OpenStudy (anonymous):

@agent0smith i know they're equivalent :P but when you substitute in x and y in as the coordinates, you get an equation in terms of a and b, which isn't conventional.

OpenStudy (agent0smith):

Yeah, it looks weird.

OpenStudy (agent0smith):

Oh i see, you listed the center as (x, y) = (-7, -6)